# Right for Right Reasons: Large Language Models for Verifiable Commonsense Knowledge Graph Question Answering

### v1

@@ Abstract
Knowledge Graph Question Answering (KGQA) methods seek to answer Natural Language questions using the relational information stored in Knowledge Graphs (KGs). With the recent advancements of Large Language Models (LLMs) and their remarkable reasoning abilities, there is a growing trend to leverage them for KGQA. However, existing methodologies have only focused on answering factual questions, e.g., "In which city was Silvio Berlusconi’s first wife born?", leaving questions involving commonsense reasoning that real-world users may pose more often, e.g., "Do I need separate visas to see the Venus of Willendorf and attend the Olympics this summer?" unaddressed. In this work, we first observe that existing LLM-based methods for KGQA struggle with hallucination on such questions, especially on queries targeting long-tail entities (e.g., non-mainstream and recent entities), thus hindering their applicability in real-world applications especially since their reasoning processes are not easily verifiable. In response, we propose Right for Right Reasons ($R^{3}$), a commonsense KGQA methodology that allows for a verifiable reasoning procedure by axiomatically surfacing intrinsic commonsense knowledge of LLMs and grounding every factual reasoning step on KG triples. Through experimental evaluations across three different tasks—question answering, claim verification, and preference matching—our findings showcase $R^{3}$ as a superior approach, outperforming existing methodologies and notably reducing instances of hallucination and reasoning errors.
@@ 1 Introduction
Commonsense reasoning is one of the most significant capabilities offered by LLMs Shen and Kejriwal (2023a). Therefore, it may seem straightforward to leverage the LLMs to reason over a set of retrieved KG facts to perform commonsense KGQA. However, LLMs are still susceptible to introducing ungrounded or incorrect information to their reasoning process – a phenomenon called hallucination Ye et al. (2023); Tonmoy et al. (2024). In conducting commonsense KGQA, LLMs may exhibit hallucinations both by introducing ungrounded factual information as well as making incorrect commonsense inferences. Hence, verifiability of the reasoning process is crucial to ensure the reliability of the final answer, especially in high-stake applications. Regrettably, none of the existing LLM-enhanced KGQA methodologies answer queries following a verifiable scheme.
@@ 2.1 Reasoning with Large Language Models
Despite being originally designed for text generation, LLMs have shown outstanding performance when applied to several other NLP sub-fields Chang et al. (2023). Particularly, the reasoning capability of LLMs has attracted considerable interest in AI research Arora et al. (2022); Sun et al. (2022). Several works have studied different reasoning skills of LLMs such as arithmetic reasoning Yuan et al. (2023), logical reasoning Liu et al. (2023), and commonsense reasoning Bian et al. (2023); Shen and Kejriwal (2023b). These abilities make LLMs apt candidates for being used as a reasoner in tasks from other domains Ren et al. (2023); Song et al. (2023); Clusmann et al. (2023).
@@ 2.2 Commonsense Question Answering
The general knowledge and conception about the world that humans possess, and their ability to reason about it is called commonsense reasoning and is a crucial cognitive ability of humans. It is also an important reasoning skill based on which AI agents are evaluated Liu et al. (2021); bauer-etal-2022-nlp; Wang et al. (2023). LLMs have shown outstanding commonsense reasoning skills and have narrowed the gap between them and humans on available datasets Guan et al. (2023a); Bian et al. (2023). Most existing datasets such as CommonsenseQA Talmor et al. (2018) and PhysicalQA Bisk et al. (2020) contain questions about concepts rather than entities. Recently, StrategyQA Geva et al. (2021a) and Creak Onoe et al. (2021b) have been proposed as datasets for commonsense reasoning about entities that can be used to introduce commonsense reasoning to KGQA.
@@ 2.3 Knowledge Graph Query Answering
The overall objective of KGQA is to answer an NL query using the KG facts. Most existing works on KGQA focus on converting the NL query into a structured formal query in a language such as SPARQL, executing the query to retrieve the required knowledge, and finally reasoning over the retrieved facts to obtain the final answer. This idea, referred to as semantic parsing Lan et al. (2021); Gu and Su (2022); Cheng et al. (2022), often involves the data and computationally expensive process of fine-tuning with thousands of labeled examples Chen et al. (2021); Shu et al. (2022). Recently, KB-BINDER has suggested a training-free semantic parsing methodology using the in-context learning ability of LLMs with few-shot examples Li et al. (2023b). Novel LLM-based methods beyond semantic parsing approach have also been proposed. KAPING Baek et al. (2023) introduced an efficient LLM-enhanced KGQA model that finds the relevant sub-graph to the query via dense retrieval and uses the LLM to reason over it in a zero-shot manner. KGR Guan et al. (2023b) proposed the idea of allowing LLMs to make claims, retrofitting those claims on the KG facts, and finally reasoning using the corrected claims. However, all existing works on KGQA are designed to answer factoid queries, and none of them has considered queries involving commonsense reasoning.
@@ 3.1 Problem Formulation
In this paper, we propose a methodology for performing commonsense KGQA that is easily extended to other related tasks such as claim verification and KG-based preference matching. The input to the problem is a NL sentence posed by the user that can be either a question in the form of an interrogative sentence, or a claim or need expressed as an imperative sentence. We use the term query, denoted by $q$, to refer to the input in all cases. The query mentions a set of anchor entities $\mathcal{E}^{q}$. A KG $K=(\mathcal{E},\mathcal{R})$ is assumed to be given, where $\mathcal{E}$ and $\mathcal{R}$ denote its set of entities and relations respectively, such that $\mathcal{E}^{q}\subset\mathcal{E}$. The objective is to follow a sequence of reasoning steps $\mathcal{S}^{q}$ to find $a^{q}\in\mathcal{O}^{q}$, the answer to the query, such that verifying the correctness of every $s^{q}_{i}\in\mathcal{S}^{q}$ is possible. Here, $\mathcal{O}^{q}$ denotes the set of possible options.
Figure 1: An example workflow of commonsense KGQA procedure using $R^{3}$.
@@ 3.2 Right for Right Reasons
Our proposed methodology casts the problem of commonsense KGQA as a tree-structured search, in which every reasoning step is either grounded on KG facts, or based on surfaced commonsense axioms, a key property that makes the reasoning procedure completely verifiable. The overall workflow of $R^{3}$ for answering a query is shown in Figure 1. In brief, $R^{3}$ first identifies the anchor entities of a query and obtains the relevant sub-graph for these entities. Next, it surfaces a commonsense axiom from the LLM that will guide the reasoning steps in that branch of the search tree. Then, at each depth level of the tree, it checks whether the commonsense axiom can be satisfied with the available KG facts, and if possible, provides an answer grounded on a subset of them. If the available KG triples are insufficient, by backward-chaining from the axiom, it selects the next entity to obtain its relevant KG sub-graph to continue the search. Each branch can continue up to a maximum depth, and if an answer is not obtained at its bottom, a new commonsense axiom will be surfaced which will guide search in a new branch until the search tree reaches its maximum breadth. Here, we explain each component of $R^{3}$ in more detail.
@@ 3.2.3 Sub-graph Pruning
$\begin{split}&\mathtt{SGP}({I}^{q}_{i},\mathcal{K}^{q})=\\ &\mathtt{top-k}_{t\in\mathcal{K}_{q}}(sim(\boldsymbol{t},\boldsymbol{I}^{q}_{i}))\cup\mathtt{LLM}_{T}(\mathcal{K}^{q},{I}^{q}_{i}),\\ &\mathcal{T}^{q}_{i}=\mathtt{SGP}({I}^{q}_{i},\mathcal{K}^{q}),\end{split}$ (4)
in which $sim$ denotes the Euclidean similarity between $\boldsymbol{t}$ and $\boldsymbol{I}^{q}_{i}$, the embedding vectors of the triple $t$ and the axiom $I^{q}_{i}$, $\mathtt{top-}k$ operator returns the first $k$ elements of the sorted list of triples by their similarity score in descending order, and $\mathtt{LLM}_{T}$ is an LLM-based module that returns a subset of $\mathcal{K}^{q}$ that are relevant to $I^{q}_{i}$.
@@ 3.2.5 Missing Evidence Identification
$\begin{split}e^{m}&=\texttt{MEI}(q,I^{q}_{i},\mathcal{T}_{i}),\\ \mathcal{E}^{q}_{j+1}&=\mathcal{E}^{q}_{j}\cup\{e^{m}\},\\ \mathcal{K}^{q}_{j+1}&=\mathcal{K}^{q}_{j}\cup\{(h,r,t)|(h,r,t)\in\mathcal{K}\land h\in e^{m}\},\\ \mathcal{T}_{i}&=\mathcal{T}_{i}\cup\mathtt{SGP}(\mathcal{K}^{q}_{j+1},I^{q}_{i}),\\ \end{split}$ (6)
@@ 4 Experiments
We empirically evaluate $R^{3}$ on three tasks: Question answering, claim verification, and KG-based preference matching. All tasks are closely related to KGQA and involve commonsense reasoning.
@@ 4.1 Task Description
Question Answering. In this task, a question that is formed around one or multiple KG entities and requires commonsense reasoning is provided. The reasoner is required to find the answer, which is either "Yes" or No.
Claim Verification Claim verification is very similar to question answering. Here, an imperative sentence making a claim about a number of entities is provided. The reasoner has to use the KG information to decide whether the claim is "Correct" or "Incorrect".
KG-based Preference Matching In this task, a NL query has been provided that explains the user’s requirement. Also, a personal KG containing evidence about the user’s preferences and restrictions is presented to the reasoner. The reasoner has to choose the item that matches both the user’s query and their personal restrictions.
@@ 4.2 Datasets
Question Answering Early KGQA datasets consisted of simple questions that can be answered using a single KG triple. Recently, datasets containing more complex questions by introducing multi-hop reasoning have been proposed Berant et al. (2013b); Trivedi et al. (2017); Gu et al. (2021). However, all KGQA datasets contain factoid questions, which do not require commonsense reasoning to answer. Some datasets exclusively focus on evaluating commonsense reasoning Talmor et al. (2018); Boratko et al. (2020); Sap et al. (2019), but their questions target concepts (e.g., river, mountain, etc.) rather than KG entities (e.g., specific people, locations, etc.).
@@
Claim Verification For the task of KG-based claim verification, we use Creak Onoe et al. (2021a), a dataset containing True/False claims written by crowd workers using Wikipedia. We follow a similar procedure applied to the Question Answering dataset to select 150 claims and write their long-tail counterparts.
@@
Exemplar Original Question
Did Alan Turing suffer from the same fate as Abraham Lincoln?
Exemplar Modified Question
Did Ivan Shuisky suffer from the same fate as Benny Frey?
@@
Exemplar Original Claim
The Bugs Bunny cartoons were influenced by the cartoon Rick and Morty.
Exemplar Modified Claim
Giovanni Battista Casti’s works may be influenced by Maria Grazia Lenisa’s poems.
@@
Exemplar Query
Sam: I like eating pulled meats, but not beef or chicken.
Original Options
@@
Personal KG
(Sam, occupation, painter), (Sam, age, 29), (Sam, medical condition, allium allergy), …
@@ 4.3 Experimental Setup
We compare $R^{3}$ against LLM baselines with Chain-of-Thought (CoT) prompting, both in zero-shot Kojima et al. (2022) and few-shot ($k=2$) settings  Wei et al. (2022) to evaluate the need for a KG to answering these queries, and three recent LLM-enhanced KGQA models, KB-BINDER Li et al. (2023a), KGR Guan et al. (2023b), and KAPING Baek et al. (2023). For question answering and claim verification tasks, we evaluate all methods on both original queries (targeting famous entities) and modified queries (targeting long-tail entities) to study their robustness to popularity shift. We use GPT-3.5 Turbo as the LLM for all models. In addition to accuracy, we perform human evaluation to measure factual and reasoning faithfulness. In particular, we use FActScore Min et al. (2023), which measures the percentage of atomic facts in an LLM’s response that is supported by a knowledge base, and Reasoning score, which measures the proportion of LLM responses in which there are no logical reasoning errors. For KG-based preference matching, our human evaluation consists of measuring Accuracy of Reasons which is simply the fraction of correct answers that were obtained from correct reasons.
@@ 4.4.1 Question Answering
The results for the question answering task are presented in table 2. $R^{3}$ beats all baselines, achieving an accuracy of 0.82 and 0.69 in the original and long-tail settings respectively. Although the strongest baseline, KAPING, achieves comparable accuracy, human evaluation reveals that KAPING’s answers are far less reliable than those of $R^{3}$. KB-BINDER’s performance is much lower than other methods, because KB-BINDER is a semantic parsing method that only supports factoid queries and not ones that require commonsense reasoning. Although 0-shot and few-shot CoT achieve 0.70 accuracy on the original queries, their accuracies drop significantly in the long-tail setting to 0.32 and 0.43 respectively. We also observe in the long-tail setting a sharp increase in the number of questions for which the LLM responds “I don’t know.”

### v2

@@ Abstract
Knowledge Graph Question Answering (KGQA) methods seek to answer Natural Language questions using the relational information stored in Knowledge Graphs (KGs). With the recent advancements of Large Language Models (LLMs) and their remarkable reasoning abilities, there is a growing trend to leverage them for KGQA. However, existing methodologies have only focused on answering factual questions, e.g., “In which city was Silvio Berlusconi’s first wife born?”, leaving questions involving commonsense reasoning that real-world users may pose more often, e.g., “Do I need separate visas to see the Venus of Willendorf and attend the Olympics this summer?” unaddressed. In this work, we first observe that existing LLM-based methods for KGQA struggle with hallucination on such questions, especially on queries targeting long-tail entities (e.g., non-mainstream and recent entities), thus hindering their applicability in real-world applications especially since their reasoning processes are not easily verifiable. In response, we propose Right for Right Reasons ($R^{3}$), a commonsense KGQA methodology that allows for a verifiable reasoning procedure by axiomatically surfacing intrinsic commonsense knowledge of LLMs and grounding every factual reasoning step on KG triples. Through experimental evaluations across three different tasks—question answering, claim verification, and preference matching—our findings showcase $R^{3}$ as a superior approach, outperforming existing methodologies and notably reducing instances of hallucination and reasoning errors.
Right for Right Reasons: Large Language Models for Verifiable Commonsense Knowledge Graph Question Answering
Armin Toroghi1, Willis Guo1, Mohammad Mahdi Abdollah Pour1, Scott Sanner1,2 1University of Toronto, Canada 2Vector Institute of Artificial Intelligence, Toronto, Canada {armin.toroghi, gwillis.guo, m.abdollahpour}@mail.utoronto.ca ssanner@mie.utoronto.ca
@@ 1 Introduction
Commonsense reasoning is one of the most significant capabilities offered by LLMs Shen and Kejriwal (2021); Zhao et al. (2024). Therefore, it may seem straightforward to leverage the LLMs to reason over a set of retrieved KG facts to perform commonsense KGQA. However, LLMs are still susceptible to introducing ungrounded or incorrect information to their reasoning process – a phenomenon called hallucination Ye et al. (2023); Tonmoy et al. (2024). In conducting commonsense KGQA, LLMs may exhibit hallucinations both by introducing ungrounded factual information as well as making incorrect commonsense inferences. Hence, verifiability of the reasoning process is crucial to ensure the reliability of the final answer, especially in high-stakes applications. Regrettably, none of the existing LLM-enhanced KGQA methodologies answer queries following a verifiable scheme.
@@ 2.1 Reasoning with Large Language Models
Despite being originally designed for text generation, LLMs have shown outstanding performance when applied to several other NLP sub-fields Chang et al. (2023). Particularly, the reasoning capability of LLMs has attracted considerable interest in AI research Arora et al. (2022); Sun et al. (2022); Xu et al. (2023). Several works have studied different reasoning skills of LLMs such as arithmetic reasoning Yuan et al. (2023), logical reasoning Liu et al. (2023), and commonsense reasoning Bian et al. (2023); Shen and Kejriwal (2023). These abilities make LLMs apt candidates for being used as a reasoner in specialized tasks Ren et al. (2023); Song et al. (2023); Clusmann et al. (2023).
@@ 2.2 Commonsense Question Answering
The general knowledge and conception about the world that humans possess, and their ability to reason about it is called commonsense reasoning and is a crucial cognitive ability of humans. It is also an important reasoning skill based on which AI agents are evaluated Liu et al. (2021); bauer-etal-2022-nlp; Wang et al. (2023). LLMs have shown outstanding commonsense reasoning skills and the gap between their performance and humans on available datasets has narrowed substantially Guan et al. (2023a); Bian et al. (2023). Most of these datasets such as CommonsenseQA Talmor et al. (2018) and PhysicalQA Bisk et al. (2020) contain questions about concepts rather than entities. Recently, StrategyQA Geva et al. (2021a) and Creak Onoe et al. (2021b) have been proposed as datasets for commonsense reasoning about entities that can be used to introduce commonsense reasoning to KGQA.
@@ 2.3 Knowledge Graph Query Answering
Answering questions using the relational information of KGs has recently gained significant attention Wang et al. (2024); Toroghi and Sanner (2024), with its applications ranging from healthcare Guo et al. (2022) to recommendation Toroghi et al. (2023). Most existing works on the task of answering NL queries using the KG facts, known as KGQA, focus on converting the NL query into a structured formal query in a language such as SPARQL, executing the query to retrieve the required knowledge, and finally reasoning over the retrieved facts to obtain the final answer. This idea, referred to as semantic parsing Lan et al. (2021); Gu and Su (2022); Cheng et al. (2022), often involves the data and computationally expensive process of fine-tuning with thousands of labeled examples Chen et al. (2021); Shu et al. (2022). Recently, KB-BINDER has suggested a training-free semantic parsing methodology using the in-context learning ability of LLMs with few-shot examples Li et al. (2023b). Novel LLM-based methods beyond semantic parsing approach have also been proposed. KAPING Baek et al. (2023) introduced an efficient LLM-enhanced KGQA model that finds the relevant sub-graph to the query via dense retrieval and uses the LLM to reason over it in a zero-shot manner. KGR Guan et al. (2023b) proposed the idea of allowing LLMs to make claims, retrofitting those claims on the KG facts, and finally reasoning using the corrected claims. However, all existing works on KGQA are designed to answer factoid queries, and none of them has considered queries involving commonsense reasoning.
@@ 3.1 Problem Formulation
In this paper, we propose a methodology for performing commonsense KGQA that is easily extended to other related tasks such as KG-based preference matching. The input to the problem is a NL sentence posed by the user that can be either a question in the form of an interrogative sentence, or a claim or need expressed as an imperative sentence. We use the term query, denoted by $q$, to refer to the input in all cases. The query mentions a set of anchor entities $\mathcal{E}^{q}$. A KG $K=(\mathcal{E},\mathcal{R})$ is assumed to be given, where $\mathcal{E}$ and $\mathcal{R}$ denote its set of entities and relations respectively, such that $\mathcal{E}^{q}\subset\mathcal{E}$. The objective is to follow a sequence of reasoning steps $\mathcal{S}^{q}$ to find $a^{q}\in\mathcal{O}^{q}$, the answer to the query, such that verifying the correctness of every $s^{q}_{i}\in\mathcal{S}^{q}$ is possible. Here, $\mathcal{O}^{q}$ denotes the set of possible options.
Figure 1: Workflow of commonsense KGQA procedure using $R^{3}$. After extracting entities from the query and retrieving their relevant sub-graphs, a commonsense axiom is surfaced from the LLM that guides the reasoning branch. After pruning sub-graphs to obtain their relevant facts to the axiom, an iterative process using the LLM is executed to either provide a fact-grounded answer or identify missing information and retrieve it from the KG. If the answer is not found after a certain depth, a new axiom is surfaced to guide a new branch.
@@ 3.2 Right for Right Reasons
Our proposed method casts the problem of commonsense KGQA as a tree-structured search, in which every reasoning step is either grounded on KG facts, or based on surfaced commonsense axioms, a key property that makes the reasoning procedure completely verifiable. The overall workflow of $R^{3}$ for answering a query is shown in Figure 1. In brief, $R^{3}$ first identifies the anchor entities of a query and obtains the relevant sub-graph for these entities. Next, it surfaces a commonsense axiom from the LLM that will guide the reasoning steps in that branch of the search tree. Then, at each depth level of the tree, it checks whether the commonsense axiom can be satisfied with the available KG facts, and if possible, provides an answer grounded on a subset of them. If the available KG triples are insufficient, by backward-chaining from the axiom, it selects the next entity to obtain its relevant KG sub-graph to continue the search. Each branch can continue up to a maximum depth, and if an answer is not obtained at its bottom, a new commonsense axiom will be surfaced which will guide search in a new branch until the search tree reaches its maximum breadth. Components of $R^{3}$ are explained here, and a series of analyses on their roles and significance are provided in Appendix A.
@@ 3.2.3 Sub-graph Pruning
$\begin{split}&\mathtt{SGP}({I}^{q}_{i},\mathcal{K}^{q})=\\ &\;\;\;\mathtt{top\textrm{-}k}_{t\in\mathcal{K}_{q}}(sim(\boldsymbol{t},\boldsymbol{I}^{q}_{i}))\cup\mathtt{LLM}_{T}(\mathcal{K}^{q},{I}^{q}_{i}),\\ &\mathcal{T}^{q}_{i}=\mathtt{SGP}({I}^{q}_{i},\mathcal{K}^{q}),\end{split}$ (4)
in which $sim$ denotes the Euclidean similarity between $\boldsymbol{t}$ and $\boldsymbol{I}^{q}_{i}$, the embedding vectors of the triple $t$ and the axiom $I^{q}_{i}$, $\mathtt{top}\textrm{-}k$ operator returns the first $k$ elements of the sorted list of triples by their similarity score in descending order, and $\mathtt{LLM}_{T}$ is an LLM-based module that returns a subset of $\mathcal{K}^{q}$ that are relevant to $I^{q}_{i}$.
@@ 3.2.5 Missing Evidence Identification
$e^{m}$ $=\texttt{MEI}(q,I^{q}_{i},\mathcal{T}_{i}),$ (6)
$\mathcal{E}^{q}_{j+1}$ $=\mathcal{E}^{q}_{j}\cup\{e^{m}\},$
$\mathcal{K}^{q}_{j+1}$ $=\mathcal{K}^{q}_{j}\cup\{(h,r,t)|(h,r,t)\in\mathcal{K}\land h\in e^{m}\},$
$\mathcal{T}_{i}$ $=\mathcal{T}_{i}\cup\mathtt{SGP}(\mathcal{K}^{q}_{j+1},I^{q}_{i}),$
@@ 4 Experiments
We empirically evaluate $R^{3}$ on three tasks: Question answering, claim verification, and KG-based preference matching. All tasks are closely related to KGQA and involve commonsense reasoning. We release all our implementation codes and data [^1] https://anonymous.4open.science/r/RRR-4F47/.
@@ 4.1 Task Description
Question Answering. In this task, a question requiring commonsense reasoning formed around some KG entities is asked. The reasoner is required to find the answer, which is either "Yes" or No.
Claim Verification Claim verification is very similar to question answering. Here, an imperative sentence including a claim about some entities is stated. The reasoner has to use the KG facts to decide whether the claim is "Correct" or "Incorrect".
KG-based Preference Matching In this task, a query explaining the user’s preference and a personal KG containing evidence about the user’s preferences and restrictions is presented to the reasoner. The reasoner has to choose the item that matches both the user’s query and their personal restrictions.
@@ 4.2 Datasets
Question Answering Early KGQA datasets consisted of simple questions that can be answered using a single KG triple. Recently, datasets containing more complex questions by introducing multi-hop reasoning have been proposed Berant et al. (2013b); Trivedi et al. (2017); Gu et al. (2021). However, all KGQA datasets contain factoid questions, which do not require commonsense reasoning to answer Guo et al. (2024). Some datasets exclusively focus on evaluating commonsense reasoning Talmor et al. (2018); Boratko et al. (2020); Sap et al. (2019), but their questions target concepts (e.g., river, mountain, etc.) rather than KG entities (e.g., specific people, locations, etc.).
@@
Claim Verification For KG-based claim verification, we use Creak Onoe et al. (2021a), a dataset containing True/False claims written by crowd workers using Wikipedia. We follow a similar procedure applied to the QA dataset to select 150 claims and write their long-tail counterparts.
@@
Exemplar Original Question Did Alan Turing suffer from the same fate as Abraham Lincoln?
Exemplar Modified Question Did Ivan Shuisky suffer from the same fate as Benny Frey?
@@
Exemplar Original Claim The Bugs Bunny cartoons were influenced by the cartoon Rick and Morty.
Exemplar Modified Claim Giovanni Battista Casti’s works may be influenced by Maria Grazia Lenisa’s poems.
@@
Exemplar Query Sam: I like eating pulled meats, but not beef or chicken.
@@
Personal KG (Sam, occupation, painter), (Sam, age, 29), (Sam, medical condition, allium allergy), …
@@ 4.3 Experimental Setup
We compare $R^{3}$ against LLM baselines with Chain-of-Thought (CoT) prompting, both in zero-shot Kojima et al. (2022) and few-shot ($k=2$) settings  Wei et al. (2022) to evaluate the need for a KG to answer these queries, and three recent LLM-based KGQA models, KB-BINDER Li et al. (2023a), KGR Guan et al. (2023b), and KAPING Baek et al. (2023). For question answering and claim verification tasks, we evaluate all methods on both original queries (targeting famous entities) and modified queries (targeting long-tail entities) to study their robustness to popularity shift. We use GPT-3.5 Turbo as the LLM for all models. In addition to accuracy, we perform human evaluation to measure factual and reasoning faithfulness. In particular, we use FActScore Min et al. (2023), which measures the percentage of atomic facts in an LLM’s response supported by a knowledge base, and Reasoning score, which measures the proportion of LLM responses in which there are no logical reasoning errors. For preference matching, our human evaluation consists of measuring Accuracy of Reasons which is the fraction of correct answers that were obtained from correct reasons.
@@ 4.4.1 Question Answering
The results for the question answering task are presented in table 2. $R^{3}$ beats all baselines, achieving an accuracy of 0.82 and 0.73 in the original and long-tail settings respectively. Although the strongest baseline, KAPING, achieves comparable accuracy, human evaluation reveals that KAPING’s answers are far less reliable than those of $R^{3}$. KB-BINDER’s performance is much lower than other methods, because KB-BINDER is a semantic parsing method that only supports factoid queries and not ones that require commonsense reasoning. Although 0-shot and few-shot CoT achieve 0.70 accuracy on the original queries, their accuracies drop significantly in the long-tail setting to 0.32 and 0.43 respectively. We also observe in the long-tail setting a sharp increase in the number of questions for which the LLM responds “I don’t know.”